OpenStudy (anonymous):

Solve both, using the substitution method. 2x + 4y = 16 x – 7 = – 4y y + 2x = 7 14 – 4x = 2y

OpenStudy (anonymous):

thank you but could you write the steps that way i could do it the next time

OpenStudy (shamil98):

For 2x + 4y = 16 Simplify this to (divide all terms by 2) x + 2y = 8 x = 7 - 4y Substitute for x and solve for y now.

OpenStudy (anonymous):

thanks!!!